OpenStudy (anonymous):

I cant figure out how to derive this: f(x)= cos((1+x^2)/(1-x^2)) And then putting in f(sqrt(2)) Thank you!

OpenStudy (anonymous):

add the number

OpenStudy (anonymous):

what i meant was f ' (sqrt(2))

OpenStudy (anonymous):

but i have trouble deriving this

ganeshie8 (ganeshie8):

use chain rule

ganeshie8 (ganeshie8):

\(\large f(x) = \cos\left(\dfrac{1+x^2}{1-x^2}\right)\) \(\large f'(x) = -\sin\left(\dfrac{1+x^2}{1-x^2}\right) \times \left(\dfrac{1+x^2}{1-x^2}\right) '\)

ganeshie8 (ganeshie8):

^differentiate the inside funciton again - use quotient/product rule

OpenStudy (anonymous):

ooh ofcourse! thank you

ganeshie8 (ganeshie8):

Once you're done, do verify ur answer against wolfram's : http://www.wolframalpha.com/input/?i=%5Ccos%5Cleft%28%5Cdfrac%7B1%2Bx%5E2%7D%7B1-x%5E2%7D%5Cright%29%27

ganeshie8 (ganeshie8):

u wlc :)

OpenStudy (anonymous):

thank you soo much! i totally get it now!

ganeshie8 (ganeshie8):

np :) good to hear :D